OpenStudy (anonymous):

Find the foci,F1 and F2 , of the hyperbola x^2-4y^2-4x-8y-16=0

OpenStudy (anonymous):

the formula of the hyperbola is : (x-2)^2/19 -( 2Y+1)^2/19 =1 => F1(-\[\sqrt{38}\] ;-1/2) F2( sqrt 38 ; -1/2)

OpenStudy (anonymous):

sorry it should be (x-2)^2 /16 - (2y+2)^2 /16 =1

OpenStudy (anonymous):

F1(-2;-1) F2(2;-1)